Circular No. 6388

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) D. Despois, Observatoire de Bordeaux; J. Wink, R. Neri, R. Lucas,

M. Grewing, S. Guilloteau, and A. Dutrey, Institut de Radioastronomie Millimétrique (IRAM), Grenoble; and P. Colom, D. Bockelée-Morvan, E. Gérard, H. Rauer, N. Biver and J. Crovisier, Observatoire de Paris-Meudon, report: “We have clearly detected the two main hyperfine components, and marginally seen the third, of the HCN (1-0) transition at 88.6 GHz in comet C/1996 B2 on Mar. 23.1 UT at the IRAM Plateau de Bure Interferometer (four 15-m antennas). The synthesized beam was 3′′.8 × 2′′.4 (baseline range 50–200 m). The emission source is compact (3′′ apparent size from a Gaussian fit) and located within ∼ 1′′ of the predicted optical ephemeris position. Averaging the data circularly around the fitted position, the velocity-integrated visibility (fringe amplitude counted in flux density units) decreases from 0.39 ± 0.05 to 0.08 ± 0.07 Jy km/s for baselines 75 to 165 m, averaged in a channel 1 km/s wide that is centered on the main hyperfine component. This is much less than the corresponding velocityintegrated single-dish flux density of 4 ± 0.1 Jy km/s in the 55′′ primary beam; the ratio seems to agree with optically-thin parent-molecule emission. The same line was again imaged on Mar. 29.6 (visibility-integrated over the central channel 1 km/s wide), yielding 0.75 ± 0.11 Jy km/s with an antenna baseline separation of 15 m, 0.19 ± 0.05 at 45 m, 0.21 ± 0.07 at 75 m, and 0.08 ± 0.04 at 165 m; the single-dish flux integrated over the same channel was 2.5 Jy km/s. The CO (2-1) line at 230 GHz was observed at the same time, with a synthesized beam of 1′′.5, yielding 3.7 ± 0.4 Jy km/s with an antenna baseline separation of 15 m, 1.22 ± 0.2 at 45 m, 0.59 ± 0.18 at 75 m, and 0.25 ± 0.13 at 165 m; the single-dish flux (22′′ primary beam) was 20 Jy km/s. For both molecules, the peak of the velocity-integrated emission is displaced by ∼ 3′′ from the computed nucleus position for Mar. 29. On both days and for both molecules, the emission peak in individual channel maps seems to shift with velocity, which may be due to jet emission.”

Total visual magnitude estimates: Apr. 18.80 UT, 2.4 (V. Znojil, Brno, Czech Republic, naked eye); 19.81, 1.4 (K. Sárneczky, Ráktanya, Hungary, naked eye; 65^{o} tail in p.a. 40^{o}); 20.14, 1.5 (R. Keen, Mt. Thorodin, CO, naked eye; 96^{o} tail in p.a. 31^{o}); 21.82, 1.9 (M. V. Zanotta, Milan, Italy, 7×42 binoculars); 22.86, 2.8 (A. Pereira, Cabo da Roca, Portugal, naked eye); 23.80, 1.8 (M. L. Paradowski, Lublin, Poland, 7×35 binoculars); 25.18, 1.8 (C. E. Spratt, Victoria, BC, 11×80 binoculars).

1996 April 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green